# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ Appendix D Proof of Lemmas
where the last inequality follows from both Lemma C.1 and C.2.
@@
Recall that a sub-Gaussian random vector $\bm{x}\in\mathbb{R}^{d}$ satisfies
@@
$=\frac{1}{n}\sum\limits_{i=1}^{n}\left(\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x};1]\rangle-\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle+\mathbb{E}\left[\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right]-\mathbb{E}\left[\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right]\right)$
@@
For the last term, we consider the event $\bigcap\limits_{i=1}^{n}\mathcal{F}_{i}$. Under this event it holds with probability at least $1-n^{-11}$ that
@@
$\widehat{\bm{M}}_{2}-\bm{M}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})-\mathbb{E}\left[\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})\right]+\frac{1}{n}\sum_{i=1}^{n}z_{i}(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})$
@@
$\quad\leq\mathbb{E}\left[\frac{\alpha}{\alpha+\theta}\cdot\exp\left\{\frac{X^{\alpha^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}\left({\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}\vee 1\right)}\right\}+\frac{\theta}{\alpha+\theta}\cdot\exp\left\{\frac{Y^{\theta^{-1}}}{\left({\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}\vee 1\right){\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\right]$
@@
Let $\{X_{i}\}_{i=1}^{n_{x}}$ be sub-Gaussian random variables and $\{Y_{i}\}_{i=1}^{n_{y}}$ be sub-exponential random variables, then we have
@@
Next, for some $l,m\in[d]$, let the Kronecker delta $\delta_{lm}=1$ only when $l=m$ and 00 otherwise. Then we have for all $i\in[n]$ that
@@
[…]
@@
where $\operatorname{VC-dim}(\mathcal{D})$ denotes the Vapnik-Chervonenkis dimension of $\mathcal{D}$. Recall that each elements of $\mathcal{P}_{k,d}(\varphi)$ is given as the intersection of $k$ jointly $s$-sparse halfspaces. Since the VC-dim of a single halfspace in $\mathbb{R}^{s}$ is $s+1$ [Csikos et al., 2018, Theorem A], we have
$\sup_{\bm{x}_{1},\dots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{P}_{k,d}(\varphi),\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{s+1}\right)^{k(s+1)},\quad\forall\varphi\in\mathcal{Z}_{s}.$ (119)
Furthemore, since $\mathcal{P}_{k,d,s}=\underset{\varphi\in\mathcal{Z}_{s}}{\bigcup}\mathcal{P}_{k,d}(\varphi)$, it follows that
$\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}\}_{i=1}^{n})=\underset{\varphi\in\mathcal{Z}_{s}}{\bigcup}\mathcal{H}(\mathcal{P}_{k,d}(\varphi),\{\bm{x}_{i}\}_{i=1}^{n}).$ (120)
@@
$\sup\limits_{\begin{subarray}{c}C\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in C\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|$
$\quad\leq\sup_{\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}}~{}\sup_{\{\omega_{i}\}_{i=1}^{n}\in\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}^{\prime}\}_{i=1}^{n})}\underbrace{\sup_{\varphi\in\mathcal{Z}_{s}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|}_{f(\omega_{1},\dots,\omega_{n})}.$ (122)
@@
D.5 Proof of Lemma A.1

### v2

@@ Appendix D Proof of Lemmas
where the last inequality follows from Lemma C.2.
@@
In what follows, we provide a proof for a stronger result that holds under a weaker assumption that $\bm{x}$ is a sub-Gaussian random vector symmetric about the origin, i.e. $\bm{x}\overset{\text{d}}{\sim}-\bm{x}$. Recall that a sub-Gaussian random vector $\bm{x}\in\mathbb{R}^{d}$ satisfies
@@
$=\frac{1}{n}\sum\limits_{i=1}^{n}(\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x};1]\rangle-\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle$
$\quad+\mathbb{E}\left[\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right]-\mathbb{E}\left[\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right])$
@@
Let $\{r_{i}\}_{i=1}^{n}$ be independent copies of a Rademacher random variable. Notice that $\bm{x}_{i}\overset{\text{d}}{\sim}\bm{x}_{i}r_{i}$ for every $i\in[n]$ by the symmetry of the Gaussian distribution. Therefore under the event $\bigcap\limits_{i=1}^{n}\mathcal{F}_{i}$, it holds that
$\mathbb{P}\left({\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{5\sigma_{z}\sqrt{\log n}}\bm{x}_{i}\right\rVert}_{\infty}\geq t\right)$ $=\mathbb{P}\left({\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{5\sigma_{z}\sqrt{\log n}}\bm{x}_{i}r_{i}\right\rVert}_{\infty}\geq t\right)$
$\leq 2\mathbb{P}\left({\left\lVert\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}r_{i}\right\rVert}_{\infty}\geq t\right)$
$=2\mathbb{P}\left({\left\lVert\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\right\rVert}_{\infty}\geq t\right),$
where the first inequality follows from the contraction principle in [Ledoux and Talagrand, 2013, Theorem 4.4]. Therefore, under the event $\bigcap\limits_{i=1}^{n}\mathcal{F}_{i}$, it holds with probability at least $1-2n^{-11}$ that
@@
$\widehat{\bm{M}}_{2}-\bm{M}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})-\mathbb{E}\left[\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})\right]$
$\quad+\frac{1}{n}\sum_{i=1}^{n}z_{i}(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})$
@@
Let $\{X_{i}\}_{i=1}^{n_{x}}$ and $\{Y_{i}\}_{i=1}^{n_{y}}$ be collections of sub-Gaussian and sub-exponential random variables respectively, then we have
@@
Next, for some $l,m\in[d]$, let the Kronecker delta $\delta_{lm}=1$ only when $l=m$ and 0 otherwise. Then we have for all $i\in[n]$ that
@@
$\leq\frac{1}{\pi_{j}^{\star}}\sum_{j^{\prime}\neq j}\mathbb{P}\left(\langle[\bm{x};\ 1],\bm{\theta}_{j^{\prime}}\rangle\geq\langle[\bm{x};\ 1],\bm{\theta}_{j}\rangle,\langle[\bm{x};\ 1],\bm{\theta}_{j}^{\star}\rangle\geq\langle[\bm{x};\ 1],\bm{\theta}_{j^{\prime}}^{\star}\rangle\right)$
$\leq\frac{1}{\pi_{j}^{\star}}\sum_{j^{\prime}\neq j}\mathbb{P}\left(\langle[\bm{x};\ 1],\bm{v}_{j,j^{\prime}}\rangle\langle[\bm{x};\ 1],\bm{v}_{j,j^{\prime}}^{\star}\rangle\leq 0\right)$
@@
where $\operatorname{VC-dim}(\mathcal{D})$ denotes the Vapnik-Chervonenkis dimension of $\mathcal{D}$. Recall that each elements of $\mathcal{P}_{k,d}(\mathcal{U})$ is given as the intersection of $k$ jointly $s$-sparse halfspaces. Since the VC-dim of a single halfspace in $\mathbb{R}^{s}$ is $s+1$ [Csikos et al., 2018, Theorem A], we have
$\sup_{\bm{x}_{1},\dots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{P}_{k,d}(\mathcal{U}),\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{s+1}\right)^{k(s+1)},\quad\forall\mathcal{U}\in\mathcal{Z}_{s}.$ (131)
Furthemore, since $\mathcal{P}_{k,d,s}=\underset{\mathcal{U}\in\mathcal{Z}_{s}}{\bigcup}\mathcal{P}_{k,d}(\mathcal{U})$, it follows that
$\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}\}_{i=1}^{n})=\underset{\mathcal{U}\in\mathcal{Z}_{s}}{\bigcup}\mathcal{H}(\mathcal{P}_{k,d}(\mathcal{U}),\{\bm{x}_{i}\}_{i=1}^{n}).$ (132)
@@
$\sup\limits_{\begin{subarray}{c}C\in\mathcal{P}_{k,d,s}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in C\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\mathcal{U}}\right\|$
$\quad\leq\sup_{\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}}~\sup_{\{\omega_{i}\}_{i=1}^{n}\in\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}^{\prime}\}_{i=1}^{n})}\underbrace{\sup_{\mathcal{U}\in\mathcal{Z}_{s}}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}}\left(\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\mathcal{U}}\right\|}_{f(\omega_{1},\dots,\omega_{n})}.$ (134)
@@
D.5 Proof of Lemma B.3
By the definition of $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ in (14), we have for all $j^{\prime}\neq j\in[k]$ that
$\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta})\cap\mathcal{C}_{j^{\prime}}(\bm{\theta}^{\star})$ $\Rightarrow\langle\bm{\xi}_{i},\bm{\theta}_{j}\rangle\geq\langle\bm{\xi}_{i},\bm{\theta}_{j^{\prime}}\rangle,\langle\bm{\xi}_{i},\bm{\theta}^{\star}_{j^{\prime}}\rangle\geq\langle\bm{\xi}_{i},\bm{\theta}^{\star}_{j}\rangle$
$\Leftrightarrow\langle\bm{\xi}_{i},\bm{v}_{j,j^{\prime}}\rangle\geq 0,\langle\bm{\xi}_{i},\bm{v}^{\star}_{j,j^{\prime}}\rangle\leq 0$
$\Rightarrow\langle\bm{\xi}_{i},\bm{v}_{j,j^{\prime}}\rangle\langle\bm{\xi}_{i},\bm{v}^{\star}_{j,j^{\prime}}\rangle\leq 0$
$\Rightarrow\langle\bm{\xi}_{i},\bm{v}^{\star}_{j,j^{\prime}}\rangle^{2}\leq\langle\bm{\xi}_{i},\bm{v}_{j,j^{\prime}}-\bm{v}^{\star}_{j,j^{\prime}}\rangle^{2},$
where the last statement holds since $ab<0$ implies $|b|\leq|a-b|$ for all $a,b\in\mathbb{R}$. From [Kim and Lee, 2023, Lemma 7.4], we have that every $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ has $(\bm{v}_{j,j^{\prime}},\bm{v}_{j,j^{\prime}}^{\star})$ in
$\mathcal{M}\triangleq\left\{(\bm{v},\bm{v}^{\star}):\|\bm{v}-\bm{v}^{\star}\|\leq 2\rho\|(\bm{v}^{\star})_{1:d}\|\right\}.$ (137)
Also, define the set
$\mathcal{S}_{\bm{v},\bm{v}^{\star}}\triangleq\left\{\bm{\xi}:\langle\bm{\xi},\bm{v}^{\star}\rangle^{2}\leq\langle\bm{\xi},\bm{v}-\bm{v}^{\star}\rangle^{2}\right\}$ (138)
Therefore it suffices to show that with probability $1-\delta$, we have
$\sup_{(\bm{v},\bm{v}^{\star})\in\mathcal{M}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{\xi}_{i}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\}}\leq\frac{\pi^{2(1+\zeta^{-1})}_{\min}}{k^{2}},$ (139)
when the sample complexity satisfies (84). For fixed $\mathcal{U}\in\mathcal{Z}_{s}$, we have that $\mathcal{S}_{\bm{v},\bm{v}^{\star}}\in\mathcal{P}_{2,d}(\mathcal{U})$ and similar to (131) we have
$\sup_{\bm{x}_{1},\dots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{P}_{2,d}(\mathcal{U}),\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{s+1}\right)^{2(s+1)}.$ (140)
Therefore, for fixed $\mathcal{U}\in\mathcal{Z}_{s}$ we have by [Kim and Lee, 2023, Lemma 6.3] with probability at least $1-\delta$ that
$\sup_{(\bm{v},\bm{v}^{\star})\in\mathcal{M}}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{\xi}_{i}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\}}-\mathbb{P}\left(\bm{\xi}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\right)\right|\leq C\sqrt{\frac{\log(1/\delta)+s\log(n/s)}{n}}.$ (141)
Furthermore, this statement holds for all $\mathcal{U}\in\mathcal{Z}_{s}$ using the union bound when we inflate the probability of error by $|\mathcal{Z}_{s}|$, i.e.
$\sup_{\begin{subarray}{c}(\bm{v},\bm{v}^{\star})\in\mathcal{M}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{\xi}_{i}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\}}-\mathbb{P}\left(\bm{\xi}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\right)\right|\leq C\sqrt{\frac{\log(1/\delta)+s\log(n\vee d/s)}{n}},$ (142)
with probability at least $1-\delta$. Furthermore, by [Kim and Lee, 2023, Lemma 7.5] similar to (125), we have
$\sup_{\begin{subarray}{c}(\bm{v},\bm{v}^{\star})\in\mathcal{M}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\mathbb{P}\left(\bm{\xi}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\right)\leq C\left[(2\rho)^{2}\log(1/\rho)\right]^{\zeta}\leq C\left[\frac{R^{2}\pi_{\min}^{2\zeta^{-1}(1+\zeta^{-1})}}{k^{2}\zeta^{-1}}\right]^{\zeta}\leq C\frac{R^{2\zeta}\pi_{\min}^{2(1+\zeta^{-1})}}{k^{2}}$ (143)
Finally, choosing the numerical constant in (84) large enough and $R>0$ small enough so that
$\mathbb{P}\left[\sup_{\begin{subarray}{c}(\bm{v},\bm{v}^{\star})\in\mathcal{M}\\ \mathcal{U}\in\mathcal{Z}_{s}\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{\xi}_{i}\in\mathcal{S}_{\bm{v},\bm{v}^{\star}}\}}>\left(\frac{\pi_{\min}^{(1+\zeta^{-1})}}{k}\right)^{2}\right]\leq\delta,$ (144)
which concludes the proof.
D.6 Proof of Lemma A.1